8. Use induction (on n) to show that: (a) (2n)! > 2 (n!)?, for n > 0. (w) (%) = (2+1), for os ms m. (0) Ž -»* (*) = (–1(m
Answer #1

(a). (2n)! > 2 (n!!, for nyo For n=0, (2.0)! = 1 2010!) = 1 + (2n! ? 2In he is true For no Suppose the statement is true 7om induction on as it is true for all nem. (c). To E-04 ( h) = 4, home for osman. For ma neno . - (7) = ( °(3) - | Holno) =